OpenStudy (anonymous):

Don't understand how to simplified this: frac{ \sqrt{x^2+1}+x }{ \sqrt{x^2+1} (the bracket is suppose between the second one) Answer: \[= \frac{ \sqrt{x+\sqrt{x^2+1}} }{ 2\sqrt{x^2+1} }\]

OpenStudy (anonymous):

it mean this: not sure to the simplify this into that answer: \[y'= \frac{ 1 }{ 2\sqrt{x+\sqrt{x^2+1}} } \left[ \frac{ \sqrt{x^2+1}+x }{ \sqrt{x^2+1} } \right]\] Sorry~ (it was hard for me type it up) \[= \frac{ \sqrt{x+\sqrt{x^2+1}} }{ 2\sqrt{x^2+1} }\]

OpenStudy (rational):

You have \[y'= \frac{ 1 }{ 2\color{blue}{\sqrt{x+\sqrt{x^2+1}}} } \left[ \frac{ \sqrt{x^2+1}+x }{ \sqrt{x^2+1} } \right]\] Multiply top and bottom by \(\color{blue}{\sqrt{x+\sqrt{x^2+1}}}\), what do you get ?

OpenStudy (anonymous):

I'm not sure how to do that...

OpenStudy (rational):

do you know how to rationalize the denominator of \(\large \dfrac{a}{\sqrt{b}}\) ?

OpenStudy (anonymous):

yoy multiply top and bottom to get rid of the square root

OpenStudy (anonymous):

but I'm don't know what to do with double square roots

OpenStudy (rational):

we use the same idea here

OpenStudy (rational):

\[\begin{align}y'&= \frac{ 1 }{ 2\color{blue}{\sqrt{x+\sqrt{x^2+1}}} } \left[ \frac{ \sqrt{x^2+1}+x }{ \sqrt{x^2+1} } \right]\\~\\ &= \frac{ 1 }{ 2\color{blue}{\sqrt{x+\sqrt{x^2+1}}} } \left[ \frac{ \sqrt{x^2+1}+x }{\sqrt{x^2+1} } \right] \times\dfrac{\color{blue}{\sqrt{x+\sqrt{x^2+1}}} }{\color{blue}{\sqrt{x+\sqrt{x^2+1}}} } \\~\\ &= \frac{ 1 }{ 2 } \left[ \frac{ \sqrt{x^2+1}+x }{\sqrt{x^2+1} } \right] \times\dfrac{\color{blue}{\sqrt{x+\sqrt{x^2+1}}} }{\left(\color{blue}{\sqrt{x+\sqrt{x^2+1}}} \right)^2} \\~\\ \end{align}\]

OpenStudy (rational):

square and sqrt cancel each other : \[\require{cancel}\begin{align}y'&= \frac{ 1 }{ 2\color{blue}{\sqrt{x+\sqrt{x^2+1}}} } \left[ \frac{ \sqrt{x^2+1}+x }{ \sqrt{x^2+1} } \right]\\~\\ &= \frac{ 1 }{ 2\color{blue}{\sqrt{x+\sqrt{x^2+1}}} } \left[ \frac{ \sqrt{x^2+1}+x }{\sqrt{x^2+1} } \right] \times\dfrac{\color{blue}{\sqrt{x+\sqrt{x^2+1}}} }{\color{blue}{\sqrt{x+\sqrt{x^2+1}}} } \\~\\ &= \frac{ 1 }{ 2 } \left[ \frac{ \sqrt{x^2+1}+x }{\sqrt{x^2+1} } \right] \times\dfrac{\color{blue}{\sqrt{x+\sqrt{x^2+1}}} }{\left(\color{blue}{\sqrt{x+\sqrt{x^2+1}}} \right)^2} \\~\\ &= \frac{ 1 }{ 2 } \left[ \frac{ \sqrt{x^2+1}+x }{\sqrt{x^2+1} } \right] \times\dfrac{\color{blue}{\sqrt{x+\sqrt{x^2+1}}} }{\color{blue}{x+\sqrt{x^2+1}}} \\~\\ &= \frac{ 1 }{ 2 } \left[ \frac{ \cancel{\sqrt{x^2+1}+x }}{\sqrt{x^2+1} } \right] \times\dfrac{\color{blue}{\sqrt{x+\sqrt{x^2+1}}} }{\color{blue}{\cancel{x+\sqrt{x^2+1}}}} \\~\\ \end{align}\]

OpenStudy (anonymous):

Wow! Thank you! Did not that you can reduce that. I tried changing the square roots with exponents. silly me..

OpenStudy (anonymous):

where did you get the square on the bottom denominator?

OpenStudy (anonymous):

nevermind , i got it

OpenStudy (anonymous):

i see what you did that! Thanks!

OpenStudy (rational):

\[\large \color{blue}{\sqrt{\heartsuit}} \times \color{blue}{\sqrt{\heartsuit}} = (\color{blue}{\sqrt{\heartsuit}} )^2\]

OpenStudy (anonymous):

Thank you! :D